OpenStudy (itsmichelle29):

Help plzzzz i medal and fan The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (anonymous):

what r the answer choices I got f(x)=4(4)x , it would make sense to represent a growth 3 times per year with f(x)=4(4)3x because the exponent x is going to be 3 times grater, since you triple the yearly growth. (I think)

OpenStudy (anonymous):

@itsmichelle29

OpenStudy (anonymous):

reply if I helped :)

OpenStudy (itsmichelle29):

so whats the correct answer

OpenStudy (itsmichelle29):

f(x) = 4(4)x; growth rate 400% f(x) = 4(4)3x, growth rate 4% f(x) = 4(1.59)3x; growth rate 59% f(x) = 4(1.59)x; growth rate 4%

OpenStudy (itsmichelle29):

@_greatmath7

OpenStudy (anonymous):

I think it is B

OpenStudy (anonymous):

Yeah it is B

OpenStudy (itsmichelle29):

... sure ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good luck! ;)

OpenStudy (anonymous):

did you get it right @itsmichelle29

OpenStudy (itsmichelle29):

I havent submited it .. im doin

OpenStudy (anonymous):

oh

OpenStudy (itsmichelle29):

there problems

OpenStudy (itsmichelle29):

OTHER*

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I thought that was your last problem:)

OpenStudy (anonymous):

if you need any more help I will be glad to assist :)

OpenStudy (itsmichelle29):

Okay

OpenStudy (anonymous):

:)